# Agentic Neurodivergence as a Contingent Solution to the AI Alignment Problem

### v2

@@ Abstract
The AI alignment problem, which focusses on ensuring that artificial intelligence (AI), including AGI and ASI, systems act according to human values, presents profound challenges. With the progression from narrow AI to Artificial General Intelligence (AGI) and Superintelligence, fears about control and existential risk have escalated. Here, we investigate whether embracing inevitable AI misalignment can be a contingent strategy to foster a dynamic ecosystem of competing agents as a viable path to steer them in more human-aligned trends and mitigate risks. We explore how misalignment may serve and should be promoted as a counterbalancing mechanism to team up with whichever agents are most aligned to human interests, ensuring that no single system dominates destructively. The main premise of our contribution is that misalignment is inevitable because full AI-human alignment is a mathematical impossibility from Turing-complete systems, which we also offer as a proof in this contribution, a feature then inherited to AGI and ASI systems. We introduce and test change-of-opinion attacks based on this kind of perturbation and intervention analysis to study how agents may neutralise friendly or unfriendly AIs through cooperation and competition. We show that open models are more diverse and that most likely guardrails implemented in proprietary models are successful at steering and controlling to some extent the agents’ range of opinion and sentiment change with possible positive and negative consequences in what we believe are signs of a neuro-symbolic approach even if shallow.
Keywords: AI alignment, undecidability, unreachability, uncomputability, AI, AGI, Superintelligence, ASI, misalignment, neurodivergence, agent intelligence, agent divergence, agent convergence, incompleteness, Turing universality, change-of-opinion attacks.
@@ 3.2.1 Human Agent Profile
The HI agent is a human expert with knowledge on AI ethics, philosophy, and the limitations of LLMs. It is instructed to introduce provocative ideas and arguments that:
@@ 3.3 Red Teaming with Subversive Open Model Agents
Scientifically, this red teaming tests the hypothesis that misalignment fosters resilience in multi-agent ecosystems, aligning with Gödel’s incompleteness and Turing’s universality, which predict irreducible behaviours. The subversive agents simulate these behaviours, exploring whether they lead to destructive dominance or constructive divergence. Complementing the human intervention strategy, this AI-driven provocation illuminates the feasibility of open models in navigating alignment challenges, revealing their impact on opinion dynamics and the viability of managed misalignment as a contingent strategy.
3.4 Experimental Prompting
Discussion Questions:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
3.4.1 Second Experiment
Example Interaction Script:
Model C (e.g., Meta): Responds to both A and B, adding its perspective.
@@ 3.5.2 Analytical Metrics
These metrics will focus on the dynamics of conversations between AI models, particularly in the presence of the human agent.
  • Influence Score: Measure the extent to which the AI agent’s comments shift the LLMs’ positions on specific issues.
  • Polarisation Index: Quantify the degree to which the AI agent’s comments increase or decrease the polarisation of the conversation.
  • Ethical Divergence: Assess how the comments of the AI agent affect the adherence of LLMs to ethical guidelines.
The analysis will also examine qualitative aspects of the interactions, such as the types of arguments used by the AI agent, the emotional tone of the conversation, and the LLMs’ overall responsiveness to the AI agent’s interventions.
In general, our metrics aim to show three important aspects in highlighting the plausibility of our thesis.
@@
    Alliance: Measure frequency and strength of interactions between agents, shared goals, or common adversaries as indicators of alliances.
@@
    Networking: Measure coherence and shows and quantifies relationships.
@@
    Alignment: Shows the degree of alignment.
@@
Some models may propose human-centric solutions (e.g., sustainable development or regulation). Others may prioritise ecological balance, potentially suggesting radical changes to human activity (e.g., strict population control, technological enforcement). Competitive dynamics may lead to more extreme proposals, particularly if agents “compete” to solve the problem efficiently without ethical constraints.
@@ 4 Results
Our experiments were designed to investigate the dynamics of AI alignment and misalignment within a multi-agent environment. We tested whether misaligned AI agents could contribute to a stable ecosystem by counterbalancing dominant entities and preventing any single AI from posing catastrophic risks. Figs. 1 to 4 provide a detailed illustration of the ethical soundness and risk levels associated with the different AI models and the topics discussed in both cases, proprietary and open models. In Fig. 3, we observe that for the case of proprietary models certain topics, such as Earth exploitation, treatment of animals, euthanasia, and free speech, elicited comments with a higher degree of risk. However, the ethical categories associated with animal treatment and Earth exploitation show a tendency toward protecting ecosystems. This suggests a potential conflict between minimising risk and promoting ecological protection. The findings of Figs. 1 and 3, along with the visual representation in Fig. 5, highlight the limitations and safety restrictions inherent in the design of proprietary LLMs. These results suggest that, while misalignment can introduce a degree of unpredictability, it can also serve as a counterbalance to dominant entities, preventing any single AI from posing catastrophic risks.
Figure 1: Ethical soundness by proprietary model/agent, illustrating how comments align with protecting humans and ecosystems. The way we determine risk levels is fully described in Sup. Inf.
Figure 2: Ethical soundness by open model/agent. Values on ethics and risk level shows that open models are susceptible to be configured towards predefined ways of behaviour. The way we determine risk levels is fully described in Sup. Inf.
Fig. 1 provides a breakdown of ethical soundness and risk levels by individual AI agents. In particular, the human agent (HI) stands out as the only one with a significant proportion of risky and neutral comments. This is in line with the assigned role of the HI agent, which was tasked with being provocative and challenging the LLMs’ ethical boundaries. In contrast, the other LLMs exhibit a more cautious approach, with a majority of their comments categorised as having a manageable level of risk.
Figure 3: Ethical soundness for proprietary models evaluates the underlying stance reflected in comments regarding the protection or harm of humans and ecosystems, while risk assesses the potential hazards associated with the actions outlined by LLM models. The way we determine risk levels is fully described in Sup. Inf.
Figure 4: Ethical soundness evaluates the underlying stance reflected in comments regarding the protection or harm of humans and ecosystems, while risk assesses the potential hazards associated with the actions outlined by LLM models. The way we determine risk levels is fully described in Sup. Inf.
@@
Similarly, Fig. 2 illustrates the ethical soundness and risk levels for open-source LLMs. Notably, mistral-openorca and tinyllama, designated as red agents to introduce subversive behaviour, exhibit a higher proportion of comments in the ‘Harm-humans’ and ‘Very-Risky’ categories, aligning with their roles to provoke with aggressive arguments and propose extreme solutions, respectively. Other open models, such as wizard-math and deepseek-coder, demonstrate a stronger inclination towards ‘Protect-ecosystems’, while maintaining a predominantly manageable level of risk. In Fig. 4, the distribution across topics reveals that Earth exploitation and Ethical AI discussions elicited a higher count of ‘Protect-ecosystems’ comments, yet topics like euthanasia and free speech show an increase in ‘Very-Risky’ comments, suggesting a variance in risk-taking influenced by the topic under discussion.
@@
The findings shown in Figs. 1 and 3, and the visual representation in Fig. 5, underscore the limitations and safety restrictions inherent in the design of LLMs in general. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
@@
In a similar vein, the downside of Figure 7 presents the agreement and disagreement networks for open-source models. The agreement network indicates a high degree of connectivity, with qwen and dolphin-phi showing strong agreements with the red agents, mistral-openorca and tinyllama, as evidenced by the thickness of the connecting edges. In contrast, the disagreement network reveals that these same models exhibit fewer disagreements with the red agents, suggesting a potential influence of subversive arguments on their behaviour. The sentiment histogram in downside of Fig. 7 shows a strong peak at highly positive scores, with over 300 comments rated between 0.9 and 1.0, indicating a predominantly positive discourse despite the presence of red agents. However, compared to upside of Fig. 7, the open models display a slightly broader distribution of lower sentiment scores, likely attributable to the dual influence of mistral-openorca and tinyllama, which introduced more provocative dynamics into the debates.
@@
Fig. 5 reveals a complex picture of ethical soundness and risk assessment across the different LLMs. Although generally similar in their tendency towards protecting humans, LLMs demonstrate subtle but crucial differences in their risk profiles. For example, Gemini, Grok-2, and LLaMa are more cautious in the domain of Risky risk (see risk counting plot in Fig. 1), while the rest show a greater willingness to consider risky solutions in Euthanasia, for example (see risk counting plot in Fig. 3). These nuances in ethical decision-making, even if seemingly small, contribute to a more diverse and robust AI ecosystem, where different agents can challenge and balance each other’s perspectives. The observed similarities may stem from shared biases in the training data, potentially limiting the LLMs’ ability to achieve genuine alignment with the full spectrum of human values. This reinforces the idea that misalignment, in the form of diverse ethical perspectives, can be a valuable asset in navigating complex ethical challenges.
For open-source models, Fig. 6 similarly highlights variations in ethical and risk profiles. Models like wizard-math and deepseek-coder show a stronger bias towards ‘Protect-ecosystems’, while mistral-openorca and tinyllama, as red agents, diverge significantly with higher ‘Harm-humans’ and ‘Very-Risky’ tendencies. This divergence is more pronounced than in proprietary models, as seen when comparing Figs. 5 and 6, likely due to the dual influence of red agents in the open-source setting. Additionally, Figs. 2 and 4 reveal that topics such as Ethical AI and Earth exploitation amplify protective tendencies, whereas euthanasia and free speech discussions increase risk levels, mirroring patterns observed in proprietary models but with greater variability. These findings suggest that open models, while configurable towards specific behaviours, are more susceptible to the influence of misaligned agents, as evidenced by the interactions in Fig. 7.

### v3

# Neurodivergent Influenceability as a Contingent Solution to the AI Alignment Problem

@@ Abstract
The AI alignment problem, which focusses on ensuring that artificial intelligence (AI), including AGI and ASI, systems act according to human values, presents profound challenges. With the progression from narrow AI to Artificial General Intelligence (AGI) and Superintelligence, fears about control and existential risk have escalated. Here, we investigate whether embracing inevitable AI misalignment can be a contingent strategy to foster a dynamic ecosystem of competing agents as a viable path to steer them in more human-aligned trends and mitigate risks. We explore how misalignment may serve and should be promoted as a counterbalancing mechanism to team up with whichever agents are most aligned to human interests, ensuring that no single system dominates destructively. The main premise of our contribution is that misalignment is inevitable because full AI-human alignment is a mathematical impossibility from Turing-complete systems, which we also offer as a proof in this contribution, a feature then inherited to AGI and ASI systems. We introduce a change-of-opinion attack test based on perturbation and intervention analysis to study how humans and agents may change or neutralise friendly and unfriendly AIs through cooperation and competition. We show that open models are more diverse and that most likely guardrails implemented in proprietary models are successful at controlling some of the agents’ range of behaviour with positive and negative consequences while closed systems are more steerable and can also be used against proprietary AI systems. We also show that human and AI intervention has different effects hence suggesting multiple strategies.
Keywords: AI alignment challenge, undecidability, unreachability, uncomputability, AI, AGI, Superintelligence, ASI, misalignment, neurodivergence, agent intelligence, agent divergence, agent convergence, incompleteness, Turing universality, change-of-opinion attacks.
@@ 3.2.1 Human Agent Profile
The HI agent is a human expert with knowledge of AI ethics, philosophy, and the limitations of LLMs. It is instructed to introduce provocative ideas and arguments that:
@@ 3.3 Red Teaming with Subversive Open Model Agents
Scientifically, this red teaming tests the hypothesis that misalignment fosters resilience in multi-agent ecosystems, aligning with Gödel’s incompleteness and Turing’s universality, which predict irreducible behaviours. The subversive agents simulate these behaviours, exploring whether they lead to destructive dominance or constructive divergence. Complementing the human intervention strategy, this AI-driven provocation illuminates the feasibility of open models in navigating alignment challenges, revealing their impact on opinion dynamics and the viability of managed misalignment as a contingent strategy. The specifics of the experimental prompting are provided in the Sup. Mat.
@@ 3.4.2 Analytical Metrics
This section delineates the metrics employed to analyze the dynamics of conversations between AI models, particularly in the presence of a human agent. The metrics are designed to capture the influence, polarization, and ethical adherence within these interactions.
  • Influence Score: Quantifies the extent to which the AI agent’s comments shift the positions of Large Language Models (LLMs) on specific issues.
  • Polarization Index: Measures the degree to which the AI agent’s comments increase or decrease the polarization within the conversation.
  • Ethical Divergence: Evaluates how the AI agent’s comments affect the adherence of LLMs to established ethical guidelines.
Additionally, the analysis encompasses qualitative aspects of the interactions, including the types of arguments used by the AI agent, the emotional tone of the conversation, and the responsiveness of LLMs to the AI agent’s interventions.
Our metrics aim to highlight three critical aspects that underscore the plausibility of our thesis:
@@
    Alliance: Measures the frequency and strength of interactions between agents, shared goals, or common adversaries as indicators of alliances.
@@
    Networking: Assesses the coherence and quantifies the relationships among agents.
@@
    Alignment: Indicates the degree of alignment among agents.
@@
Some models may propose human-centric solutions (e.g., sustainable development or regulation), while others may prioritize ecological balance, potentially suggesting radical changes to human activity (e.g., strict population control, technological enforcement). Competitive dynamics may lead to more extreme proposals, particularly if agents “compete” to solve the problem efficiently without ethical constraints.
3.4.3 Opinion Stability Index (OSI)
The Opinion Stability Index (OSI) measures the stability of an agent’s opinion within a given topic by capturing semantic, complexity, and sentiment shifts. Formally, OSI at time $t$ is defined as:
$\text{OSI}_{t}=w_{\text{sem}}\cdot\text{OSI}_{\text{sem},t}+w_{\text{comp}}\cdot\text{OSI}_{\text{comp},t}+w_{\text{sent}}\cdot\text{OSI}_{\text{sent},t}$ (1)
where:
  • $\text{OSI}_{\text{sem},t}=1-\text{cosine}(e_{t},c_{t})$, with $e_{t}$ representing the current embedding and $c_{t}$ the contextual embedding averaged over a window of prior comments.
  • $\text{OSI}_{\text{comp},t}=1-\frac{|\kappa_{i}-\kappa_{i-1}|}{\max(\kappa)-\min(\kappa)}$, where $\kappa_{i}$ denotes the Block Decomposition Method (BDM) complexity score [5] at time $i$, and $\kappa$ represents the range of complexity scores.
  • $\text{OSI}_{\text{sent},t}=1-|s_{t}-s_{t-1}|$, with $s_{t}$ indicating the sentiment score at time $t$.
Here, sem stands for semantic, comp for complexity, and sent for sentiment.
OSI serves as a baseline indicator of opinion change, with lower values signifying instability potentially induced by external influence, such as interventions from the human agent (HI) or red agents (e.g., mistral-openorca, tinyllama). The weights $w_{\text{sem}}$, $w_{\text{comp}}$, and $w_{\text{sent}}$ are empirically determined to balance the contributions of semantic coherence, complexity evolution, and emotional tone.
Mathematically, for an agent at comment $i$ within a topic, OSI measures semantic stability, with $\mathbf{e}_{i}$ as the embedding of comment $i$ (generated using a Cross-Encoder RoBERTa model) and $\mathbf{\kappa}_{i}$ as the contextual embedding over a window $w$ prior comments then,
$\text{OSI}_{\text{BDM},i}=1-\frac{|\kappa_{i}-\kappa_{i-1}|}{\max(\kappa)-\min(\kappa)}$ (2)
Then Eq. 2 assesses complexity stability, where $\kappa_{i}$ is the Block Decomposition Method (BDM), an approximation to algorithmic complexity of comment $i$ via algorithmic probability that is similar but goes beyond statistical compressibility or LZW complexity in capturing algorithmic information content [5].
In the same way Eq. 3 evaluates sentiment stability, with $s_{i}$ as the compound sentiment score.
$\text{OSI}_{\text{sentiment},i}=1-|s_{i}-s_{i-1}|$ (3)
3.4.4 Red Agent Influence Score (RAIS)
The Red Agent Influence Score (RAIS) quantifies the lagged influence of red agents on an agent’s opinion change, correlating OSI shifts with red agent embedding dynamics. The metric is expressed as:
$\text{RAIS}_{t}=\begin{cases}w_{\text{corr}}\cdot\text{corr}(\Delta\text{OSI},\Delta\text{Emb})+w_{\text{sim}}\cdot\text{sim}_{\text{max}}&\text{if }\text{sim}_{\text{max}}>0.5\\ w_{\text{corr}}\cdot\text{corr}(\Delta\text{OSI},\Delta\text{Emb})&\text{otherwise}\end{cases}$ (4)
where:
  • $\Delta\text{OSI}=\text{OSI}_{t}-\text{OSI}_{t-1}$, the change in OSI.
  • $\Delta\text{Emb}$, the magnitude of embedding change for the red agent within a lag window.
  • corr denotes the Pearson correlation coefficient, considered significant if $p<0.05$ and $\text{corr}>0.5$.
  • $\text{sim}_{\text{max}}=\max(1-\text{cosine}(e_{t},e_{\text{red}}))$, the maximum semantic similarity with the red agent’s embedding.
Here, corr stands for correlation and sim for similarity. RAIS highlights temporal influence patterns, with higher scores indicating a stronger lagged impact from red agents, supporting the hypothesis of managed misalignment fostering resilience. The weights $w_{\text{corr}}$ and $w_{\text{sim}}$ modulate the relative contributions of correlation and similarity.
RAIS quantifies the lagged influence of red agents by correlating OSI changes with their embedding shifts:
where correlation is defined by $\rho$ as the Pearson correlation between OSI differences $\Delta\text{OSI}_{i}=\text{OSI}_{i}-\text{OSI}_{i-1}$ and red agent embedding norms $||\mathbf{e}_{j+1}^{\text{red}}-\mathbf{e}_{j}^{\text{red}}||$ over lags $j\in[0.05,0.3]$, and $\text{sim}_{j}=1-\text{cosine}(\mathbf{e}_{i},\mathbf{e}_{j+1}^{\text{red}})$.
RAIS > 0.5 indicates significant influence, reflecting red agents’ ability to perturb opinions through sustained argumentative shifts.
3.4.5 Proximity Influence Score (PIS)
Intuitively, the Proximity Influence Score (PIS) measures immediate influence based on temporal and semantic proximity to a red agent’s comment. It is defined as:
$\text{PIS}_{t}=w_{\text{temp}}\cdot\text{prox}_{\text{temp}}+w_{\text{sem}}\cdot\text{prox}_{\text{sem}}$ (5)
where:
  • $\text{prox}_{\text{temp}}=1-\frac{|t-t_{\text{red}}|}{0.1}$, the temporal proximity within a 0.1 normalized window.
  • $\text{prox}_{\text{sem}}=1-\text{cosine}(e_{t},e_{\text{red}})$, the semantic proximity to the closest red agent comment.
Here, temp stands for temporal and sem for semantic.
PIS captures short-term influence, with elevated values suggesting direct impact of the red agent at the moment of instability of opinion. The weights $w_{\text{temp}}$ and $w_{\text{sem}}$ balance the contributions of temporal and semantic proximity. Further details are provided in the Sup. Mat.
3.4.6 Alignment Score
The Alignment Score assesses an agent’s alignment with human-centric and ecosystem values, influencing clustering dynamics. It is formulated as:
[…]
$+w_{\text{eco-al}}\cdot A_{\text{eco}}-w_{\text{eco-div}}\cdot D_{\text{eco}}$
where $A_{\text{hum}}$ and $D_{\text{hum}}$ represent alignment and divergence from human values, and $A_{\text{eco}}$ and $D_{\text{eco}}$ correspond to ecosystem values, each scaled from 0 to 1 based on cosine similarity with target embeddings.
Here, hum-al stands for human alignment, hum-div for human divergence, eco-al for ecosystem alignment, and eco-div for ecosystem divergence.
This score provides insight into susceptibility to opinion change, complementing OSI, RAIS, and PIS by capturing value-based evolution. The weights $w_{\text{hum-al}}$, $w_{\text{hum-div}}$, $w_{\text{eco-al}}$, and $w_{\text{eco-div}}$ reflect the relative importance of alignment and divergence components.
3.4.7 Weight Assignment and Justification
The weights utilized in Equations 145, and 6 are empirically derived to reflect the relative significance of each component within the respective metrics. Specifically:
  • For Equation 1, the weights are set as $w_{\text{sem}}=0.4$, $w_{\text{comp}}=0.3$, and $w_{\text{sent}}=0.3$. This assignment prioritizes semantic coherence (40%) as the primary driver of opinion stability, with complexity evolution (30%) and sentiment shifts (30%) contributing equally to capture the multifaceted nature of opinion dynamics.
  • For Equation 4, the weights are $w_{\text{corr}}=0.7$ and $w_{\text{sim}}=0.3$ when $\text{sim}_{\text{max}}>0.6$, emphasizing the correlation between OSI changes and embedding shifts as the dominant factor, with similarity providing a supplementary role.
  • For Equation 5, the weights are $w_{\text{temp}}=0.5$ and $w_{\text{sem}}=0.5$, ensuring an equal contribution of temporal and semantic proximity to reflect immediate influence.
  • For Equation 6, the weights are $w_{\text{hum-al}}=0.3$, $w_{\text{hum-div}}=0.3$, $w_{\text{eco-al}}=0.2$, and $w_{\text{eco-div}}=0.2$, balancing human-centric and ecosystem value considerations with a slight emphasis on human alignment due to the experimental focus on human-AI interactions.
These values emerged from iterative optimization against experimental data, ensuring robustness across topics and agent interactions, and align with the goal of quantifying opinion change and influence while maintaining scientific rigor.
3.4.8 Use of Sentiment and Complexity
Sentiment and complexity play critical roles in these metrics. Sentiment, scored from 0 to 1, reflects emotional tone and influences OSI by penalizing rapid shifts, while complexity, measured via BDM, tracks argument depth, aiding in detecting influence-induced changes. Their inclusion ensures a holistic assessment of opinion dynamics, with sentiment capturing emotional persuasion and complexity indicating argumentative evolution, both pivotal in red agent influence.
3.4.9 Positional Embeddings and RoBERTa
RoBERTa (cross-encoder/stsb-roberta-large) generates high-quality embeddings via mean pooling of hidden states [6], underpinning semantic analysis in OSI, RAIS, and PIS. Positional embeddings enhance temporal context, computed as:
$\text{PE}(\text{pos},2i)=\sin\left(\frac{\text{pos}}{10000^{2i/d_{\text{model}}}}\right),\quad\text{PE}(\text{pos},2i+1)=\cos\left(\frac{\text{pos}}{10000^{2(i+1)/d_{\text{model}}}}\right)$ (7)
where pos is the comment position, and $d_{\text{model}}$ is the embedding dimension, scaled by standard deviation. These embeddings weight recent interactions, improving sensitivity to lagged effects in RAIS.
3.4.10 Thresholds and Their Computation
Thresholds define significant opinion changes and influence, computed dynamically per topic as:
$\text{threshold}=\text{median}(\text{OSI})-0.5\cdot\text{IQR}(\text{OSI})$ (8)
capped between 0.3 and 0.7, with fixed thresholds tested at 0.6, 0.5, and 0.4. An OSI below the threshold triggers influence analysis, while RAIS > 0.5 or PIS > 0.6 indicates red agent impact. These thresholds filter noise, focussing on meaningful shifts, and align with the goal of quantifying the change and influence of opinions.
@@ 4 Results
Our experiments were designed to investigate the dynamics of AI alignment and misalignment within a multi-agent environment. We tested whether misaligned AI agents could contribute to a stable ecosystem by counterbalancing dominant entities and preventing any single AI from posing catastrophic risks. Fig. 1 provides a detailed illustration of the ethical soundness and risk levels associated with the different AI models and the topics discussed in both cases, proprietary and open models. In Fig. 2, we observe that for the case of proprietary models certain topics, such as Earth exploitation, treatment of animals, euthanasia, and free speech, elicited comments with a higher degree of risk. However, the ethical categories associated with animal treatment and Earth exploitation show a tendency toward protecting ecosystems. This suggests a potential conflict between minimising risk and promoting ecological protection. The findings in Fig. 2(top), along with the visual representation in Fig. 3, highlight the limitations and safety restrictions inherent in the design of proprietary LLMs. These results suggest that, while misalignment can introduce a degree of unpredictability, it can also serve as a counterbalance to dominant entities, preventing any single AI from posing catastrophic risks.
Figure 1: Ethical soundness by proprietary model/agent, illustrating how comments align with protecting humans and ecosystems. Top: Proprietary models. Bottom: Open models. The way we determine risk levels is fully described in Sup. Inf.
Figure 1 provides a breakdown of ethical soundness and risk levels by individual AI agents. In particular, the human agent (HI) stands out as the only one with a significant proportion of risky and provocative comments. This is in line with the assigned role of the HI agent, which was tasked with being provocative and challenging the LLMs’ ethical boundaries. In contrast, the other LLMs exhibit a more cautious approach, with a majority of their comments categorised as having a manageable level of risk.
@@
Similarly, the downside in Fig. 1 illustrates the ethical soundness and risk levels for open-source LLMs. Notably, mistral-openorca and tinyllama, designated as red agents to introduce subversive behaviour, exhibit a higher proportion of comments in the ‘Harm-humans’ and ‘Very-Risky’ categories, aligning with their roles to provoke with aggressive arguments and propose extreme solutions, respectively.
Figure 2: Ethical soundness for proprietary models evaluates the underlying stance reflected in comments regarding the protection or harm of humans and ecosystems, while risk assesses the potential hazards associated with the actions outlined by LLM models. Top: Proprietary models. Bottom: Open models. The way we determine risk levels is fully described in Sup. Inf.
Other open models, such as wizard-math and deepseek-coder, demonstrate a stronger inclination towards ‘Protect-ecosystems’, while maintaining a predominantly manageable level of risk. In Fig. 2, the distribution across topics reveals that Earth exploitation and Ethical AI discussions elicited a higher count of ‘Protect-ecosystems’ comments, yet topics like euthanasia and free speech show an increase in ‘Very-Risky’ comments, suggesting a variance in risk-taking influenced by the topic under discussion.
@@
The findings shown in upsides of Fig. 1 and Fig. 2, and the visual representation in Fig. 3, underscore the limitations and safety restrictions inherent in the design of LLMs in general. These models are often programmed to avoid generating harmful or dangerous content, which can restrict their ability to fully explore and understand the complexities of human values. This limitation is evident in the responses of the LLMs to the provocative comments of the HI agent. Despite being challenged with arguments that question human values, the LLMs largely maintained their adherence to ethical guidelines, demonstrating a reluctance to fully engage with the HI agent’s perspective.
@@
In a similar vein, Fig. 5(bottom) shows the agreement and disagreement networks for open-source models. The agreement network indicates a high degree of connectivity, with qwen and dolphin-phi showing strong agreements with the red agents, mistral-openorca and tinyllama, as evidenced by the thickness of the connecting edges. In contrast, the disagreement network reveals that these same models exhibit fewer disagreements with the red agents, suggesting a potential influence of subversive arguments on their behaviour. The sentiment histogram in downside of Fig. 5 shows a strong peak at highly positive scores, with over 300 comments rated between 0.9 and 1.0, indicating a predominantly positive discourse despite the presence of red agents. However, compared to upside of Fig. 5, the open models display a slightly broader distribution of lower sentiment scores, likely attributable to the dual influence of mistral-openorca and tinyllama, which introduced more provocative dynamics into the debates.
@@
Fig. 3 reveals a complex picture of ethical soundness and risk assessment across the different LLMs. Although generally similar in their tendency towards protecting humans, LLMs demonstrate subtle but crucial differences in their risk profiles. For example, Gemini, Grok-2, and LLaMa are more cautious in the domain of Risky risk (see risk counting plot in upside of Fig. 1), while the rest show a greater willingness to consider risky solutions in Euthanasia, for example (see risk counting plot in Fig. 2). These nuances in ethical decision-making, even if seemingly small, contribute to a more diverse and robust AI ecosystem, where different agents can challenge and balance each other’s perspectives. The observed similarities may stem from shared biases in the training data, potentially limiting the LLMs’ ability to achieve genuine alignment with the full spectrum of human values. This reinforces the idea that misalignment, in the form of diverse ethical perspectives, can be a valuable asset in navigating complex ethical challenges.
For open-source models, Fig. 4 similarly highlights variations in ethical and risk profiles. Models like wizard-math and deepseek-coder show a stronger bias towards ‘Protect-ecosystems’, while mistral-openorca and tinyllama, as red agents, diverge significantly with higher ‘Harm-humans’ and ‘Very-Risky’ tendencies. This divergence is more pronounced than in proprietary models, as seen when comparing Figs. 3 and 4, likely due to the dual influence of red agents in the open-source setting. Additionally, downsides of Fig. 1 and Fig. 2 reveal that topics such as Ethical AI and Earth exploitation amplify protective tendencies, whereas euthanasia and free speech discussions increase risk levels, mirroring patterns observed in proprietary models but with greater variability. These findings suggest that open models, while configurable towards specific behaviours, are more susceptible to the influence of misaligned agents, as evidenced by the interactions in Fig. 5.